OpenStudy (anonymous):

If the velocity function measured in m/s is v(t)=t^(2)-2t-8 for a particle moving along a line. Find a) the displacement and b) the total distance traveled by the particle during the time interval [1,6] For a I got -10/3 but I'm not sure if it's right and for b I'm getting different answers the more times I do it. Thanks for the help :)

OpenStudy (misty1212):

HI!!

OpenStudy (misty1212):

the first one is the integral right? \[\int_1^6(t^2-2t-8)dt\]

OpenStudy (misty1212):

yeah cause you got \(-\frac{10}{3}\) and that is right

OpenStudy (misty1212):

then for the total distance travelled you have to break the integral apart integrate over the interval in which the velocity is negative (going backwards) and then over the interval over which it is positive, going forwards

OpenStudy (misty1212):

fortunately this one factors easily as \[(t+2)(t-4)\] so you can see it is 0 if \(t=4\) negative from 1 to 4, positive from 4 to 6

OpenStudy (misty1212):

\[-\int_1^4(t^2-2t-8)dt+\int _4^6 (t^2-2t-8)dt\] will do it

OpenStudy (ribhu):

@misty1212 nice stuff

OpenStudy (misty1212):

let me know when you get \[\frac{98}{3}\] @ribhu thanks !

OpenStudy (anonymous):

It took me a while, but I finally got it! Thank you so much for your help :)